OpenStudy (anonymous):

Medal OR fan please help Factor: x2 - 9x + 12

OpenStudy (johnweldon1993):

We can use the quadratic formula to solve this \[\large \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] You have the equation x² - 9x + 12 so your a = 1 b = -9 c = 12 Plug those into that equation and you'll find your 2 answers...

OpenStudy (anonymous):

(x + 1)(x - 12) (x - 2)(x + 6) (x - 2)(x - 6) prime I think its prime right?

OpenStudy (johnweldon1993):

Right it is prime...just wanted to make sure you knew what that formula was :) yeah because the 2 numbers in those parenthesis while they DO multiply to = 12...non add to make -9 so yes...it is prime

OpenStudy (anonymous):

okay thank you!

OpenStudy (johnweldon1993):

No problem good work!